OpenStudy (anonymous):

What are the area and perimeter of the sector of the circle shown? C is the center. Use an EXACT answer with fractions

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

okay do you know the formula for a cector? area and perimeter/?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

i have to hurry on this one 2, only have a certain time limit

OpenStudy (anonymous):

\[\theta/360 *\pi*r^2\]

OpenStudy (anonymous):

and yea i have to go so lets finish...so 90/360=0.25 0.25*pi*3^2 0.25*pi*9 simplify

OpenStudy (anonymous):

that is your area

OpenStudy (anonymous):

7.065 is area? @atreyu6s3x6

OpenStudy (anonymous):

is that using the exact answer in fractions? @atreyu6s3x6

OpenStudy (anonymous):

A=7.068 without roundong

OpenStudy (anonymous):

@MyNameIsNemo can you help her?

OpenStudy (anonymous):

im sorry i have to go now............. i have class now... D: im sorry

OpenStudy (mynameisnemo):

uhm maybe possibly? .-.

OpenStudy (anonymous):

@MyNameIsNemo we found area now we need perimeter

OpenStudy (mynameisnemo):

meh okie i can try

OpenStudy (anonymous):

okay thanks @MyNameIsNemo

OpenStudy (mynameisnemo):

idk the equation for the perimeter .-. @AlexandervonHumboldt2 might help you though, hes very smart and helpful see if you can get him

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

we found area, now we need perimeter

OpenStudy (alexandervonhumboldt2):

for P use formula 2*pi*r*C/360=2*3.14*3*90/360=4.71

OpenStudy (alexandervonhumboldt2):

where C is central angle

OpenStudy (anonymous):

so is 4.71 my answer for perimeter? @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah if they mean arc length by perimeter

OpenStudy (anonymous):

yes, can you help me with one more fast question?

OpenStudy (alexandervonhumboldt2):

ok